y’ = Csc x -yCot x
1
Expert's answer
2020-11-17T17:44:05-0500

"y' = csc x - ycotx\\
y' = \dfrac{1}{sinx} - y\dfrac{cosx}{sinx}\\
y = u(x)*v(x)\\
y' = u'v+uv'\\
u'v+uv'+uv\dfrac{cosx}{sinx} = \dfrac{1}{sinx}\\
\space\\
u'v +u(v'+v\dfrac{cosx}{sinx}) =\dfrac{1}{sinx}\\
\space\\
v' +v\dfrac{cosx}{sinx}=0\\
\space\\
\dfrac{dv}{dx} =-v\dfrac{cosx}{sinx}\\
\space\\
\dfrac{dv}{v} = -cotxdx\\
\space\\
\int \dfrac{dv}{v} =- \int cotxdx\\
ln v = -ln(sinx)\\
v = \dfrac{1}{sinx} = cscx\\
u' * \dfrac{1}{sinx} = \dfrac{1}{sinx}\\
\dfrac{du}{dx} = 1\\
\space\\
du = dx\\
\space\\
\int du =\int dx\\
\space\\
u = x +C, \space C-constant\\
y = u*v =csc*(x+C) = Ccscx+xcscx\\
answer: y = Ccscx+xcscx"